OpenStudy (anonymous):

solve by substitution r+s=-12 4r-6s=12

OpenStudy (anonymous):

1. Solve for either variable in the first equation 2. Substitute that information into the second equation

OpenStudy (anonymous):

S=12-r so use this to substitute into the equation 4r-6s=12 So you would get 4r-6(12-r)=12 Distribute the 6(12-r) you get 72-6r Your equation should look like: 4r-72+6r=12 Add like terms. Now your equation is 10r-72=12 Add 72 to both sides 10r=84 r=8.4 Now fill 8.4 into the first equation to get the s. 8.4+s=-12 subtract 8.4 s=-20.4